# KCL-PH-TH-2025-49

### v1

@@ Abstract
Noncommutative gravity, based on a twist-deformation of the differential geometry of spacetime and a first-order formulation of the dynamics, requires additional gravitational degrees of freedom as well as an enlargement of the gauge group of Lorentz transformations of the tetrad frame. As such, it offers a theoretical playground to build fundamentally motivated extension to general relativity. The dynamical degrees of freedom include a ${\rm GL}(2,\mathbb{C})$ gauge connection and two independent tetrads. The theory allows for interaction terms between the two tetrads, whose structure displays some similarities with ghost-free bigravity. The extra gravitational degrees of freedom survive in the commutative limit. We show the effective action obtained in this limit, discuss its symmetries, and compare it with other bigravity theories. The dynamics of homogeneous and isotropic cosmological solutions split into two branches. One is characterized by a constant and purely spatial curvature two-form. The other displays a richer gauge freedom, and the Hamiltonian analysis of the dynamics reveals three extra first-class constraints in addition to the generator of time reparametrizations.
@@ 2 Noncommutative gravity with bimetric interactions
Here, the $c_{n}$ are free coupling constants and the dynamical variables are represented by the bitetrad $\bf e$} and the ${\rm GL}(2,\mathbb{C})$ gauge connection $\boldsymbol{\omega}$}, respectively defined as [^1] Note the imaginary prefactor of $\tilde{e}^{I}$ in the definition (2.2a), unlike Refs. [8, 7] where the prefactor is real. This redefinition ensures that both $e^{I}$ and $\tilde{e}^{I}$ contribute positively to the kinetic term.
@@
${\boldsymbol{\omega}}$ $=\frac{1}{2}\omega^{IJ}\Gamma_{IJ}+\omega\mathds{1}+\bar{\omega}\gamma_{5}~,$ (2.2b)
@@
and its Hodge dual (2.1) is defined as ${\bf F}=-i{\bf F}\gamma_{5}$ . $\bf F$} can be decomposed onto the generators of the Lie algebra of ${\rm GL}(2,\mathbb{C})$ as
@@
By construction, the action (2.1) is invariant under diffeomorphisms and under $\star$-deformed ${\rm GL}(2,\mathbb{C})$ gauge transformations
@@ 3 Commutative limit and comparison with bimetric gravity theories
where $P_{IJKL}\coloneqq\frac{1}{2}\varepsilon_{IJKL}+\frac{2}{\beta}\eta_{I[K}\eta_{L]J}$ and $A\coloneqq-i\omega+\overline{\omega}/\beta$ is a one-form associated to a ${\rm U}(1)$ gauge symmetry. We note that the interaction terms with couplings $c_{2}$, $c_{3}$ in the action (2.1) are exactly vanishing in the commutative limit. In the following, we will refer to the action obtained in the commutative limit (3.1) as the NCG action for brevity.
Although the NCG action (3.1) is obtained in the commutative limit, it is quite different from the first-order formulation of GR. This is due to the presence of extra dynamical fields, which survive in the limit, that are associated with the enlarged symmetry ${\rm GL}(2,\mathbb{C})$. We stress that the introduction of such extra fields is necessary to ensure consistency between gauge symmetry and the underlying noncommutative structure. Non-trivial interaction terms further enrich the structure of the theory.
@@
Written in this form, the action is manifestly under $U(1)$ gauge transformations $E^{I}\to e^{-i\varphi}E^{I}$ , $\overline{E}^{I}\to e^{i\varphi}\overline{E}^{I}$. This is equivalent to the symmetry that mixes $e^{I}$ and $\tilde{e}^{I}$ which is included in (3.2).
@@ 5 Cosmological dynamics and Hamiltonian analysis
$e^{I}\wedge e^{J}+\tilde{e}^{I}\wedge\tilde{e}^{J}=2n\left(a^{2}+b^{2}c\right)\delta^{[I}_{0}\delta^{J]}_{i}\mbox{d}t\wedge dx^{i}+\left(a^{2}+b^{2}\right)\delta^{[I}_{i}\delta^{J]}_{j}\mbox{d}x^{i}\wedge\mbox{d}x^{j}~.$ (5.4)
Interestingly, solutions to the connection equation (4.1) fall into two different branches, depending on whether (5.4) is purely spatial or has non-zero time-space components (equivalently, whether $a^{2}+b^{2}c$ vanishes or not). The two branches are analyzed separately in the following sections.
@@ 5.1 Branch I
On this branch $a^{2}+b^{2}c\neq 0$ , and the solution to the connection equation (4.1) reads
@@
Given that $h$ is real, from Eq. (5.8b) we conclude that, on this branch $c_{1}<0$ is necessary to ensure consistency of the dynamics. The general solution to the system (5.8) can be given in terms of three free functions of time, namely $u(t)$, $v(t)$, and $w(t)$ :
@@
with $\mathcal{C}_{1}\coloneqq\sqrt{-\tilde{c}_{1}}/M_{\rm Pl}$ . The fact that the general solution is completely expressed in terms of free functions is a peculiar property of the system, which depends on the high degree of gauge symmetries. A detailed comparison with the dynamics in the Alexandrox-Speziale is carried out in the Appendix A. In Section 6, we further investigate the dynamical structure of the theory in its symmetry-reduced cosmological sector, to shed light on the gauge redundancies displayed by the exact solution (5.9).
@@ 5.2 Branch II
Thus, we have six unknown independent components of the spin connection $\omega^{I}_{\mu J}$ , namely $\omega^{0}_{11}$ , $\omega^{0}_{12}$ , $\omega^{0}_{13}$ , $\omega^{0}_{23}$ , $\omega^{0}_{22}$ , $\omega^{0}_{33}$ that are left to determine. Moreover, we obtain the algebraic constraint $a^{2}+b^{2}=\mathcal{C}^{2}$, with $\mathcal{C}$ an arbitrary real constant. Hence, the evolution of the two scale factors corresponds to the motion of a point particle on a circle with constant radius, with arbitrary velocity. In turn, this constraint implies that the two-form (5.4), besides being purely spatial on this branch, is also constant. The evolution of the scale factors and the relative lapse can be expressed as, $a=\mathcal{C}\cos\theta(t)$ , $b=\mathcal{C}\sin\theta(t)$ , $c=-(\tan\theta(t))^{-2}$ , with $\theta(t)$ an arbitrary function of time. This arbitrariness in the scale factors is not fixed by the remaining equations of motion, as shown below.
@@ 6.2 Canonical analysis
From here we obtain the primary Hamiltonian $H_{P}$ as the sum of $H_{C}$ and all the primary constrains $\phi_{i}$ with Lagrange multipliers $\lambda_{i}$ , that is, $H_{P}\coloneqq H_{C}+\lambda_{i}\phi_{i}$ , with $i=1,\dots,5$ . Next, we need to ensure that the primary constraints are preserved by time evolution. This is achieved by imposing the consistency conditions $\dot{\phi}_{i}=[\phi_{i},H_{P}]=0$ , [^4] The Poisson bracket is defined as usual by $F,G]=\frac{\partial F}{\partial\alpha}\frac{\partial G}{\partial\pi_{\alpha}}-\frac{\partial F}{\partial\pi_{\alpha}}\frac{\partial G}{\partial\alpha} [$32]. which lead to one secondary constraint
@@
Second-class constraints can be used to eliminate entirely some canonical variables from the description. Solving the constraints (6.11) as strong equations (i.e., restricting to the hypersurfaces they define in phase space), we can eliminate, for instance, $\pi_{b}$ and $\pi_{h}$ . At this point, the Poisson bracket must be replaced by the Dirac bracket, defined as $F,G]_{*}\coloneqq[F,G]-[F,\chi_{\alpha}]\mathcal{M}^{\alpha\beta}[\chi_{\beta},G]$ , where $\chi_{\alpha}$ is the set of second class constraints and $\mathcal{M}^{\alpha\beta}$ is the inverse of the matrix of the Poisson brackets of all the second-class constraints. In reduced phase space, the first-class constraints read
@@
Once all the gauge transformations have been identified, counting the number of physical degrees of freedom becomes straightforward. The key point is that the canonical gauge-fixing procedure requires the same number of new constraints as the number of first-class constraints in the theory (so that the gauge-fixing constraints form pairs of second-class constraints with the corresponding first-class constraints). For a satisfactory set of gauge conditions, one is left with second-class constraints only, and therefore the total Hamiltonian is uniquely determined [34]. Consequently, the initial ten-dimensional phase space undergoes a reduction by two due to the presence of the second-class constraints $\Phi_{5}$ and $\Phi_{6}$ that restrict the initial data, in addition to twice the number of first-class constraints, i.e., $2\times 4=8$ . Hence, we are led to the conclusion that the system has exactly zero physical degrees of freedom! Thus, the dynamics for this system are pure gauge.
@@ 7 Discussion
Acknowledgments
MdC acknowledges support from INFN iniziativa specifica GeoSymQFT. MS acknowledges support from the Science and Technology Facility Council (STFC), UK, under the research grant ST/X000753/1. ASO acknowledges financial support from the fellowship PIF21/237 of the UPV/EHU. This work falls within the scopes of the COST Actions CA23130 “Bridging high and low energies in search of quantum gravity” (BridgeQG), CA21136 “Addressing observational tensions in cosmology with systematics and fundamental physics” (CosmoVerse), CA23115 “Relativistic Quantum Information” (RQI).

### v2

@@ Abstract
Noncommutative gravity, based on a twist-deformation of the differential geometry of spacetime and a first-order formulation of the dynamics, requires additional gravitational degrees of freedom as well as an enlargement of the gauge group of Lorentz transformations of the tetrad frame. As such, it offers a theoretical playground to build fundamentally motivated extensions to general relativity. The dynamical degrees of freedom include a ${\rm GL}(2,\mathbb{C})$ gauge connection and two independent tetrads. The theory allows for interaction terms between the two tetrads, whose structure displays some similarities with ghost-free bigravity. The extra gravitational degrees of freedom survive in the commutative limit. We show the effective action obtained in this limit, discuss its symmetries, and compare it with other bigravity theories. The dynamics of homogeneous and isotropic cosmological solutions split into two branches. One is characterized by a constant and purely spatial curvature two-form. The other displays a richer gauge freedom, and the Hamiltonian analysis of the dynamics reveals three extra first-class constraints in addition to the generator of time reparametrizations.
@@ 2 Noncommutative gravity with bimetric interactions
Here, $\wedge_{\star}$ denotes the deformed wedge product [6, 8], whose properties are reviewed below in this section. The $c_{n}$ are free coupling constants and the dynamical variables are represented by the bitetrad $\bf e$} and the ${\rm GL}(2,\mathbb{C})$ gauge connection $\boldsymbol{\omega}$}, respectively defined as [^1] Note the imaginary prefactor of $\tilde{e}^{I}$ in the definition (2.2a), unlike Refs. [8, 7] where the prefactor is real. This redefinition ensures that, in the commutative limit, both $e^{I}$ and $\tilde{e}^{I}$ contribute with the same sign to the kinetic term in the action (3.1)..
@@
${\boldsymbol{\omega}}$ $=\frac{1}{2}\omega^{IJ}\Gamma_{IJ}+\omega\mathds{1}+\tilde{\omega}\gamma_{5}~,$ (2.2b)
@@
where $\rm d$} denotes the exterior derivative.The Hodge dual of the field strength can be expressed in this notation as ${\bf F}=-i{\bf F}\gamma_{5}$ . Moreover, $\bf F$} can be decomposed onto the generators of the Lie algebra of ${\rm GL}(2,\mathbb{C})$ as
@@
By construction, the action (2.1) is invariant under diffeomorphisms and under $\star$-deformed ${\rm GL}(2,\mathbb{C})$ gauge [8, 7]
@@ 3 Commutative limit and comparison with bimetric gravity theories
where $P_{IJKL}\coloneqq\frac{1}{2}\varepsilon_{IJKL}+\frac{2}{\beta}\eta_{I[K}\eta_{L]J}$ and $A\coloneqq-i\omega+\tilde{\omega}/\beta$ is a one-form associated to a ${\rm U}(1)$ gauge symmetry. We note that the interaction terms with couplings $c_{2}$, $c_{3}$ in the action (2.1) are exactly vanishing in the commutative limit. In the following, we will refer to the action obtained in the commutative limit (3.1) as the NCG action for brevity.
Although the NCG action (3.1) is obtained in the commutative limit, it is quite different from the first-order formulation of GR. This is due to the presence of extra dynamical fields, which survive in the limit, that are associated with the enlarged symmetry ${\rm GL}(2,\mathbb{C})$. We stress that the introduction of such extra fields is necessary to ensure consistency between gauge symmetry and the underlying noncommutative structure. Non-trivial interaction terms further enrich the structure of the theory. We observe that deviations from GR persist also in the absence of interaction terms between tetrads (i.e., in the special case $c_{1}=c_{4}=0$). Moreover, such deviations are not a peculiarity of the pure gravity theory here considered, and in fact persist also in the presence of matter couplings, unless further conditions are imposed on the dynamical fields. In particular, in Ref. [8], where fermionic couplings are included, the continuity of the $\theta\to 0$ limit is ensured by imposing a suitable charge conjugation condition on the extra dynamical fields (here denoted $\tilde{e}^{I}$, $\omega$, $\tilde{\omega}$), such that they vanish in the limit.
@@
Written in this form, the action is manifestly invariant under $U(1)$ gauge transformations $E^{I}\to e^{-i\varphi}E^{I}$ , $\overline{E}^{I}\to e^{i\varphi}\overline{E}^{I}$. This is equivalent to the symmetry that mixes $e^{I}$ and $\tilde{e}^{I}$ which is included in (3.2).
@@ 4 Field equations
Here $\mathcal{D}_{\omega}$ denotes the exterior covariant derivative,
$\mathcal{D}_{\omega}\left(e^{I}\wedge e^{J}+\tilde{e}^{I}\wedge\tilde{e}^{J}\right)={\rm d}\left(e^{I}\wedge e^{J}+\tilde{e}^{I}\wedge\tilde{e}^{J}\right)+\omega^{I}{}_{K}\wedge\left(e^{K}\wedge e^{J}+\tilde{e}^{K}\wedge\tilde{e}^{J}\right)+\omega^{J}{}_{K}\wedge\left(e^{I}\wedge e^{K}+\tilde{e}^{I}\wedge\tilde{e}^{K}\right)~.$ (4.2)
@@ 5 Cosmological dynamics and Hamiltonian analysis
$e^{I}\wedge e^{J}+\tilde{e}^{I}\wedge\tilde{e}^{J}=2n\left(a^{2}+b^{2}c\right)\delta^{[I}_{0}\delta^{J]}_{i}\mbox{d}t\wedge\mbox{d}x^{i}+\left(a^{2}+b^{2}\right)\delta^{[I}_{i}\delta^{J]}_{j}\mbox{d}x^{i}\wedge\mbox{d}x^{j}~.$ (5.4)
Taking the exterior covariant derivative of this expression and substituting it into the connection equation (4.1) gives
$\begin{split}\Big[(a\dot{a}+b\dot{b})\delta^{[I}_{i}\delta^{J]}_{j}-n(a^{2}+b^{2}c)\left(\omega^{[I}_{i}{}_{0}\,\delta^{J]}_{j}+\omega^{[I}_{i}{}_{j}\,\delta^{J]}_{0}\right)+(a^{2}+b^{2})&\omega^{[I}_{0}{}_{i}\,\delta^{J]}_{j}\Big]\mbox{d}t\wedge\mbox{d}x^{i}\wedge\mbox{d}x^{j}\\ &+(a^{2}+b^{2})\omega^{[I}_{l}{}_{i}\,\delta^{J]}_{j}\mbox{d}x^{l}\wedge\mbox{d}x^{i}\wedge\mbox{d}x^{j}=0~.\end{split}$ (5.5)
Interestingly, solutions to the connection equation (5.5) fall into two different branches, depending on whether the two-form (5.4) is purely spatial or has non-zero time-space components (equivalently, whether $a^{2}+b^{2}c$ vanishes or not). The two branches are analyzed separately in the following sections.
@@ 5.1 Branch I
On this branch $a^{2}+b^{2}c\neq 0$ , and then, from (5.5), and using the antisymmetry of $\omega^{IJ}$ , the solution to the connection equation (4.1) reads
@@
Given that $h$ is real, from Eq. (5.9b) we conclude that, on this branch $c_{1}<0$ is necessary to ensure consistency of the dynamics. The general solution to the system (5.9) can be given in terms of three arbitrary functions of time, namely $u(t)$, $v(t)$, and $w(t)$ :
@@
with $\mathcal{C}_{1}\coloneqq\sqrt{-\tilde{c}_{1}}/M_{\rm Pl}$ . The fact that the general solution is completely expressed in terms of free functions is a peculiar property of the system, which depends on the high degree of gauge symmetries. A detailed comparison with the dynamics in the Alexandrov-Speziale is carried out in the Appendix A. In Section 6, we further investigate the dynamical structure of the theory in its symmetry-reduced cosmological sector, to shed light on the gauge redundancies displayed by the exact solution (5.10).
@@ 5.2 Branch II
Thus, we have six unknown independent components of the spin connection $\omega^{I}_{\mu J}$ , namely $\omega^{0}_{11}$ , $\omega^{0}_{12}$ , $\omega^{0}_{13}$ , $\omega^{0}_{23}$ , $\omega^{0}_{22}$ , $\omega^{0}_{33}$ that are left to determine. Moreover, from the first equation, we obtain the algebraic constraint $a^{2}+b^{2}=\mathcal{C}^{2}$, with $\mathcal{C}$ an arbitrary real constant. Hence, the evolution of the two scale factors corresponds to the motion of a point particle on a circle with constant radius, with arbitrary velocity. In turn, this constraint implies that the two-form (5.4), besides being purely spatial on this branch, is also constant. The evolution of the scale factors and the relative lapse can be expressed as, $a=\mathcal{C}\cos\theta(t)$ , $b=\mathcal{C}\sin\theta(t)$ , $c=-(\tan\theta(t))^{-2}$ , with $\theta(t)$ an arbitrary function of time. This arbitrariness in the scale factors is not fixed by the remaining equations of motion, as shown below.
@@ 6.2 Canonical analysis
From here we obtain the primary Hamiltonian $H_{P}$ as the sum of $H_{C}$ and all the primary constrains $\phi_{i}$ with Lagrange multipliers $\lambda_{i}$ , that is, $H_{P}\coloneqq H_{C}+\lambda_{i}\phi_{i}$ , with $i=1,\dots,5$ . Next, we need to ensure that the primary constraints are preserved by time evolution. This is achieved by imposing the consistency conditions $\dot{\phi}_{i}=[\phi_{i},H_{P}]=0$ , where, along this section, [,][\,,] denotes the Poisson bracket, defined as usual by $F,G]=\frac{\partial F}{\partial\alpha}\frac{\partial G}{\partial\pi_{\alpha}}-\frac{\partial F}{\partial\pi_{\alpha}}\frac{\partial G}{\partial\alpha} [$32]. These lead to one secondary constraint
@@
Second-class constraints can be used to eliminate entirely some canonical variables from the description. Solving the constraints (6.11) as strong equations (i.e., restricting to the hypersurfaces they define in phase space), we can eliminate, for instance, $\pi_{b}$ and $\pi_{h}$ . After imposing second-class constraints strongly, Poisson brackets no longer preserve the constraints under time evolution. At this point, the Poisson bracket must be replaced by the Dirac bracket [32], defined as $F,G]_{*}\coloneqq[F,G]-[F,\chi_{\alpha}]\mathcal{M}^{\alpha\beta}[\chi_{\beta},G]$ , where $\chi_{\alpha}$ is the set of second class constraints and $\mathcal{M}^{\alpha\beta}$ is the inverse of the matrix of the Poisson brackets of all the second-class constraints. In reduced phase space, the first-class constraints read
@@
Once all the gauge transformations have been identified, counting the number of physical degrees of freedom becomes straightforward. The key point is that the canonical gauge-fixing procedure requires the same number of new constraints as the number of first-class constraints in the theory (so that the gauge-fixing constraints form pairs of second-class constraints with the corresponding first-class constraints). For a satisfactory set of gauge conditions, one is left with second-class constraints only, and therefore the total Hamiltonian is uniquely determined [34]. Consequently, the initial ten-dimensional phase space undergoes a reduction by two due to the presence of the second-class constraints $\Phi_{5}$ and $\Phi_{6}$ that restrict the initial data, in addition to twice the number of first-class constraints, i.e., $2\times 4=8$ . Hence, we are led to the conclusion that the system has exactly zero physical degrees of freedom. Thus, the dynamics for this system in the pure gravity case are pure gauge.